OpenStudy (anonymous):

Write an equivalent expression for 4* (3x) (the little asterik is a multiplication symbol) using the associative property. A. 4*(x*3) B. (3x)*4 C. (4*3)x

OpenStudy (lyrae):

associative property: a*(b*c) = (a*b)*c

OpenStudy (lyrae):

nah, look at what i did with the parentheses. See how something moved inside while somehing else moved outside...

OpenStudy (lyrae):

Quick reminder: 3x = 3*x

OpenStudy (anonymous):

Thank you so much! I appreciate your help!

OpenStudy (lyrae):

Yw, anytime :)